For the point listed below,show the inequality(s)for which it is a solution.
(5,0)
y<-3/2x+3
2x+y<10
y>-1

Thank you.

Answer by rothauserc(4718)

For point (5,0), x = 5, y = 0
:
y < -3/2x + 3
0 < -3/2(5) + 3
0 < -15/2 + 3 = -9/2 = -4.5
NO
:
2x + y < 10
2(5) + 0 < 10
10 < 10
NO
:
y > -1
0 > -1
YES
:
******************************
y > -1 is a correct solution
for the point (5,0)
******************************
: